Consider a random variable X with the following probability distribution: p(x)-0.05x. for x the the expected value of Y-6X-7 2, 3, 4, 5, or 6 Find
Answer #1

P(X = 2) = 0.05 * 2 = 0.1

P(X = 3) = 0.05 * 3 = 0.15

P(X = 4) = 0.05 * 4 = 0.2

P(X = 5) = 0.05 * 5 = 0.25

P(X = 6) = 0.05 * 6 = 0.3

E(X) = 2 * 0.1 + 3 * 0.15 + 4 * 0.2 + 5 * 0.25 + 6 * 0.3 = 4.5

E(Y) = E(6X - 7) = 6*E(X) - 7 = 6 * 4.5 - 7 = 20